) For each city across Canada, economists construct a price index for a similar basket of goods. In Edmonton (AB) the index is 132.00, and in Thunder Bay (ON) the index is 122.40. If you have been offered $130,000 for a job in Edmonton and $105,000 for a similar job in Thunder Bay, which job offers you the highest purchasing power for the bundle of goods in the price index? Use the Edmonton value as the base.


1
Expert's answer
2021-11-12T07:59:43-0500

The quantity of goods that can be bought in Edmonton with $130 000 where the price level is 132 is:

"Real\ wage =\frac{nominal\ wage}{price\ level} \\\\

Real\ wage =\frac{ \$130 000}{ 132}"


Real wage is 984.85 units of goods


The quantity of goods that can be bought in Thunder Bay with $105 000 where the price level is 122.40 is:

"Real\ wage =\frac{nominal\ wage}{price\ level} \\\\

Real\ wage =\frac{ \$105 000}{ 122.40}"

"Real\ wage = 857,84\ units\ of\ goods"


According to the estimates above, the real wage in Edmonton would be higher (based on the salary offer and current price level). This is because a salary earned in Edmonton can buy more units (or baskets) of items than an income earned in Thunder Bay at the same price level.